Simple​ Simon's Bakery purchases supplies on terms of 1.1/10, net 30. If Simple​ Simon's chooses to take the discount​ offered, it must obtain a bank loan to meet its​ short-term financing needs. A local bank has quoted Simple​ Simon's owner an interest rate of 11.2% on borrowed funds. Should Simple​ Simon's enter the loan agreement with the bank and begin taking the​ discount? ​ (Use 365 days for a​ year.)
Answer #1
Discount rate is = 1.1% and net days = 30 - 10 days = 20 Days
It means if the payment is within discount period than net payment = $ 100 - $ 1.10 = $ 98.90
Interest for discount period for early payment of 20 days ($ 100 / $ 98.90 ) - 1
Interest for discount period of 20 Days = 1.01112 -1
Interest for discount period of 20 Days = 0.01112 (Not Rounded)
Effective Annual Interest = 0.01112 X 365/20 = 0.2030
Effective Annual Interest (In percentage) = 20.30%
Annual Discount we are getting on supplies is 20.30% and banks are offering 11.2%
it means Simple Simon's can enter the loan agreement with bank because this is beneficiary for Him